OpenStudy (anonymous):

Mary has $15 and she wants to buy some earrings. She picks one pair that costs $7.99 and two pairs that cost $2.59 each. Which statement best describes if an exact total or an approximate total should be calculated?

OpenStudy (anonymous):

Where are the answer choices (statements)

OpenStudy (sptogl):

15-(8+(2.6x2))

OpenStudy (anonymous):

15-[ 7.99 + 2(2.59) ] = 15-13.17 = 1.83